# MeshVPR: Citywide Visual Place Recognition Using 3D Meshes

### v1

@@ 5.3 How 3D mesh quality affects results
Given that citywide 3D meshes play a central role in MeshVPR, we investigate how their resolution influences the robustness of the mesh-based VPR system. For this purpose, we take the two overlapping 3D mesh of San Francisco LQ and San Francisco HQ (cf. Tab. 1) and use them both to generate synthetic views of the city.
@@
Figure 4: Results with MeshVPR on High Quality (HQ) and Low Quality (LQ) meshes. Quantitative results (left) indicate a strong correlation between results and mesh quality. All results on the table are computed with MeshVPR applied to different VPR models. The gap in mesh quality is quantified in Fig. 4. Qualitative results (right) visually show how predictions are affected by the synthetically generated images. For each one of the 4 queries (i.e. the images without green/red boxes) we show the top-2 candidates with SALAD+MeshVPR on the high-quality (HQ) database (top row for each query) and the top-2 candidates with the low-quality (LQ) database.
@@
While being somewhat a predictable result, we find very promising that data quality has such a strong impact on results: in fact, as the quality (and availability) of these city-wide 3D meshes is steadily increasing, so will the results of MeshVPR.
@@ 5.4 Bridging the Syn2Real performance gap
Extensive validation presented in the above section demonstrates that MeshVPR makes each VPR algorithm obtain satisfactory performances on the mesh-based VPR setting. At this point, a natural question arises: to what extent do we recover the performance we would have obtained if a database of real images had been available?
@@ 5.7 Limitations and advantages of mesh-based VPR
Standard VPR (with real images for a database) and mesh-based VPR come each with their own strengths and weaknesses. The main strength of standard VPR is the performance that SOTA models achieve even on large-scale datasets, as shown by the results in Fig. 6, (and [3, 10, 19]). Mesh-based VPR, having to overcome the strong domain shifts between database and query images, results in lower performance even with the strong improvements given by MeshVPR. Conversely, mesh-based VPR provides a number of advantages:
@@
6 Conclusions
In this work, we propose the task of citywide mesh-based visual place recognition, which aims at obtaining scalable solutions to the problem of image localization, relying solely on a large 3D mesh. To evaluate previous VPR techniques we create three new datasets, through which we are able to quantify the performance drop of existing models on this new task. We then develop a simple yet effective technique, called MeshVPR, to align two models to the same feature space, with one model taking real images as inputs, and the other one taking synthetic images. We demonstrate how MeshVPR can be trained on any dataset providing real and synthetic images, which does not require to be from the same city as the test set. MeshVPR shows impressive results and strong flexibility, being able to train on top of any existing VPR model. In conclusion, citywide mesh-based VPR is an exciting possibility which we believe can spark a new line of research, which could unleash its full potential in allowing ubiquitous localization in practical scenarios.

### v2

@@ 5.3 How 3D mesh quality affects results
Given that citywide 3D meshes play a central role in MeshVPR, we investigate the effect of their quality on the mesh-based VPR system. For this purpose, we take the two overlapping 3D mesh of San Francisco LQ and San Francisco HQ (cf. Tab. 1) and use them both to generate synthetic views of the city (cf. Fig. 4).
@@
Albeit somewhat predictable, this result is promising for the development of mesh-based VPR: in fact, as the quality (and availability) of these city-wide 3D meshes is steadily increasing, so will the results of MeshVPR.
Figure 5: Results with MeshVPR on High Quality (HQ) and Low Quality (LQ) meshes. Quantitative results (left) indicate a strong correlation between results and mesh quality. All results on the table are computed with MeshVPR applied to different VPR models. Qualitative results (right) visually show how predictions are affected by the synthetically generated images. For each one of the 4 queries (i.e. the images without green/red boxes) we show the top-2 candidates with SALAD+MeshVPR on the high-quality (HQ) database (top row for each query) and the top-2 candidates with the low-quality (LQ) database.
@@ 5.4 Bridging the Syn2Real performance gap
Extensive validation presented in the above section demonstrates that MeshVPR makes each VPR algorithm obtain satisfactory performances on the mesh-based VPR setting. A natural question arises: to what extent do we recover the performance we would have obtained if a database of real images had been available?
@@ 5.7 Limitations and advantages of mesh-based VPR
Standard VPR (with real images for a database) and mesh-based VPR come each with their own strengths and weaknesses. The main strength of standard VPR is the performance that SOTA models achieve even on large-scale datasets, as shown by the results in Fig. 6, (and [3, 10, 19]). Mesh-based VPR, having to overcome the strong domain shifts between database and query images, results in lower performance even with the strong improvements given by MeshVPR. On the other hand, mesh-based VPR provides a number of advantages:
@@
6 Conclusions and future works
In this work, we propose the task of citywide mesh-based visual place recognition, which aims at obtaining scalable solutions to the problem of image localization, relying solely on a large 3D mesh. We create three new datasets, to quantify the performance drop of existing models on this new task. We then develop a simple yet effective technique, called MeshVPR, to align two models to the same feature space – one model taking real images as inputs, and the other one taking synthetic images. MeshVPR can be trained on any dataset of real and synthetic image pairs, on top of any existing VPR model, showing impressive results and strong flexibility.
We believe that the novel problem setup introduced in this paper opens countless future directions. For example: (1) using a full mesh-based localization pipeline based on MeshVPR plus visual localization methods as in [36, 37], (2) generating synthetic images from multiple domains (e.g. synthetic night images), (3) simulate more viewpoints (e.g. from sidewalks or parks), (4) perform drone-based localization with a synthetic set of aerial images and (5) apply style transfer to transform all the images to a single domain. In conclusion, citywide mesh-based VPR is an exciting new line of research which could unleash its full potential in allowing ubiquitous localization in practical scenarios.
Acknowledgements.
We acknowledge the Cineca award under the Iscra initiative, for the availability of high performance computing resources. This work was supported by CINI. Project supported by ESA Network of Resources Initiative. This study was carried out within the project FAIR - Future Artificial Intelligence Research - and received funding from the European Union Next-GenerationEU (Piano nazionale di ripresa e resilienza (PNRR) – missione 4 componente 2, investimento 1.3 – D.D. 1555 11/10/2022, PE00000013 - CUP: E13C22001800001). This manuscript reflects only the authors’ views and opinions, neither the European Union nor the European Commission can be considered responsible for them. European Lighthouse on Secure and Safe AI – ELSA, Horizon EU Grant ID: 101070617